OpenStudy (anonymous):

write and solve and equation to find the value of the variable. 104, 137, 154, 131, x; mean = 130

pooja195 (pooja195):

(104 + 137 + 154 + 131 + x)/5 = 130 (526 + x)/5 = 130 multiply both sides by 5: 526 + x = 650 subtract 526 from both sides: x = 124

OpenStudy (anonymous):

Thank you!

pooja195 (pooja195):

np! welcome to openstudy !!!

pooja195 (pooja195):

http://openstudy.com/code-of-conduct an no plagersm